OpenStudy (anonymous):

2x + 4y = 14 3x + 12y = 51

OpenStudy (amistre64):

if you multiply the top by -3, we can eliminate the y values and just have the xs to focus on

mathslover (mathslover):

ok : \[\large{2x+4y=14}\] Can u solve for x here ?

mathslover (mathslover):

Just take the first equation into the account ... and tell me tht wht r u getting for x ?

mathslover (mathslover):

@dezzy please reply us

OpenStudy (anonymous):

multiply first equation with 3: 6x+12y=42...........(1) 3x+12y=51...........(2) subtract 2nd equation from 1st one: 3x=-9 x=-3 put the value of x in any one of the equations and get the value of y 3(-3)+12y=51 12y=51+9 12y=60 y=5

mathslover (mathslover):

@dezzy we give live to live responses ... please tell us ..where r u having problem ..? http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Good Attempt @nitz

OpenStudy (anonymous):

thanks @lllasMcKay

OpenStudy (anonymous):

Its ILLAS MCKAY

OpenStudy (anonymous):

ok sorry.........

OpenStudy (anonymous):

thanks nitz

OpenStudy (anonymous):

you cant solve for x unless there is another equation

OpenStudy (anonymous):

Welcome spelling is wrong @mathslover

OpenStudy (anonymous):

See if there are two variables then we require two equations..

mathslover (mathslover):

\[\Huge{\mathbb{Welcome}\textbf{To}\mathbb{Open}\textbf{Study}}\]